OpenStudy (anonymous):

I COULD REALLY USE SOME HELP!!! WILL GIVE MEDALS!!!! PLEASE!!!! 1. Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, 8>, v = <-4, 8> and 2. Let u = <-4, 1>, v = <-1, 6>. Find -2u + 4v.

OpenStudy (anonymous):

psh.. nice question

OpenStudy (anonymous):

@dan815 I've seen your work i know you're good can you help me out??

OpenStudy (anonymous):

@nincompoop can you help me out??

OpenStudy (amistre64):

the usual approach for the first one is to use the dot product equivalent

OpenStudy (amistre64):

\[|u||v|\cos(a)=u\cdot v\]

OpenStudy (amistre64):

the second one is just multiplication and addition ..

OpenStudy (anonymous):

Is the first one 15.4 degrees??

OpenStudy (amistre64):

a - b = tan^(-1)(-8/5) - tan^(-1)(-8/4) im going to go with no, assuming you didnt make a typo